Valdosta State University has two colleges, College of Business Administration (COBA) & College of Education (COE). The two support departments are Physical Plant & the Library. The costs incurred in each of these departments are as follows:

COBA

$ 90,000

COE

$ 70,000

Physical Plant

$120,000

Library

$ 80,000

The actual level of support is as follows:

Supplied by

Used by

Physical Plant

Library

COBA

COE

Physical Plant

40%

35%

25%

Library

20%

50%

30%

Determine the:

  1. Cost allocated to COBA using the direct method
  2. Cost allocated to COBA using the step-down method (allocating Library first)
  3. Cost allocated to COBA using the reciprocal method
Answer #1

1. Cost allocated to COBA using Direct Method

Let the proportion of Physical Plant be ‘x’ and Library be ‘y’.

Total cost of Physical Plant (x) = $120,000 + 0.2y (20% cost of Library) ——(1)

Total cost of Library (y) = $80,000 + 0.4x (40% cost of physical plant) —— (2)

Substituting (2) in (1),

x = 120,000 + 0.2 (80,000 + 0.4x)

x = 120,000 + 16,000 + 0.08x

x - 0.08 = 136,000

0.92x =136,000

x = 147,826.09
Substituting x=$147,826.09 in (2),

y = 80,000 + 0.4 (147,826.09)

y = 139,130.44

Cost allocated to COBA = 90,000 + 35%x + 50%y

= 90,000 + 35% * 147,836.09 + 50% * 139,130.44

= $211,307.85